# Feasibility Guaranteed Learning-to-Optimize in Wireless Communication Resource Allocation

### v1

@@ VI Conclusions
This article provides a comprehensive study of feasibility-guaranteed L2O techniques in the context of wireless RA. We systematically categorized existing L2O approaches based on their constraint-handling strategies, ranging from soft penalty-based methods to hard-constrained formulations using projection and structural embeddings. To benchmark their practical applicability, we conducted experiments on a QoS-aware WSR maximization problem, revealing the superior performance of model-driven, hard-constrained methods with model-driven deep learning in balancing feasibility and optimality. Finally, we provide future opportunities to develop unique feasibility guaranteed L2O methods for wireless RA problem.

### v2

@@ VI Conclusions
This article provides a comprehensive study of feasibility-aware L2O techniques in the context of wireless RA. We systematically categorized existing L2O approaches based on their constraint-handling strategies, ranging from soft penalty-based methods to hard-constrained formulations using projection and structural embeddings. To benchmark their practical applicability, we conducted experiments on a QoS-aware WSR maximization problem, revealing the superior performance of model-driven, hard-constrained methods with model-driven deep learning in balancing feasibility and optimality. Finally, we provide future opportunities to develop unique feasibility-aware L2O methods for wireless RA problem.